Montserrat Volcano Observatory

Daily Report
Report for the period 16:00 27 December
to 16:00 28 December 1996
The current alert level is ORANGE

Activity on the October 1 dome continues at a high level with numerous rockfalls. The banded seismic tremor stopped yesterday afternoon although seismicity is still high. The changing nature of the activity is of some concern and there are dangers that it might escalate with little or no warning. An escalation could involve large pyroclastic flows into the Tar River Valley or sudden collapses from the Galway's Wall. Either could have serious consequences for much of the evacuated area. Depending on the wind direction, rockfalls from the dome may cause some ashfall in areas to the west or northwest of the volcano.

There is still no access to zone A/B. Restricted access is allowed to zone C/D, but only for essential purposes and by people with a means of rapid exit. All other zones have normal occupation.

Observations from the ground and the helicopter today showed a lot of rockfall activity on the east and north-east faces of the October 1 dome. There is a significant amount of new material on the top of the dome. This is darker in colour than the older material and is probably hotter. Many of the rockfalls produced small ash clouds which drifted to the west on the wind. There was intermittent light ashfall in the St. Patrick's and Amersham areas during much of the day.

Seismic activity has changed. The banded tremor which has dominated since 22 December died out with the last notable peak at about 11 am yesterday. Activity since then has consisted of large hybrid events and rockfall signals: 171 hybrid events and 75 rockfalls were recorded in this reporting period. Much of this activity has been randomly spaced but there have been two small peaks in the activity, at 09:00 and 14:00 today. These may represent the restarting of regular-spaced seismic activity.

Analysis of the EDM measurements made late on 26 December shows a small amount of shortening on the lines to Castle Peak, much less than the 3 cm per day recorded immediately before that. Measurements made today confirm the reduction, with 2.4 mm of shortening in two days. The drop in the deformation rate coincides approximately with the appearance of new material at the surface on the October 1 dome.

GPS measurements were made today at stations on the western side of the volcano. Theodolite measurements were taken of the heights of points on the dome to detect any recent growth. These will be used along with photographs taken today from fixed positions around the volcano.

COSPEC measurements were made yesterday and today to measure the amount of sulphur dioxide coming from the volcano. These are the first measurements since the COSPEC failed in mid-November. The data from yesterday average about 350 tonnes per day and peaked at about 400 tonnes per day shortly after one of the peaks in seismic tremor. These levels are similar to those seen during October and November. The data from today has not yet been processed.

Dr Sayadul Arafin returned to SRU today at the end of his tour of duty at MVO

Montserrat Volcano Observatory